OpenStudy (anonymous):

Write the expression logb(7/sqrtx) - logb (sqrt 7x) as a single logarithm

OpenStudy (nowhereman):

apply the logarithm rules!

OpenStudy (anonymous):

log x - log y = log (x/y) so the answer is log(7root(7x)/root(x))

OpenStudy (nowhereman):

The rule is correct, but you didn't apply it correctly. Also you can simplify that even more!

OpenStudy (anonymous):

oh sorry..division mistake. log(root(7)/x)

OpenStudy (anonymous):

Yes ive solved this 4 times and cant cat the right answer. possible answers are: \[A. logb (\sqrt{7}/x) B. logb(x ^{3/2}\sqrt{7}) C. logb(7 ^{3/2}\sqrt{x})\]

OpenStudy (anonymous):

\[D. logb \sqrt{7x} E. logb(x/\sqrt{7})\]

OpenStudy (nowhereman):

(A) is correct.